# TrackID3x3: A Dataset and Algorithm for Multi-Player Tracking with Identification and Pose Estimation in 3x3 Basketball Full-court Videos

### v1

@@ 3 TrackID3x3 Dataset
As described in Table 2, we propose TrackID3x3 dataset composed of three subsets: Indoor, Outdoor, and Drone datasets (the novelty of our dataset is described in Introduction and Table 1). In this section, we describe the data collection and annotations in this order.
3.1 Dataset collection
@@
Indoor dataset. As described in [32], Indoor dataset was constructed a fixed-camera 3x3 basketball videos consisting of 42 videos totaling 7,531 frames for six players on the court. Six female university basketball players, who regularly participate in official 5-on-5 games. The measurement was conducted in an university gymnasium (half-court of 5-on-5 basketball court) using a Sony HDR-CX680 camera ($1280\times 720$). For the mini-game design, offensive and defensive teams were fixed with players being assigned to specific roles. The ball handler of #1 is fixed, and the player begins without directly passing to #3. The start positions on the two offensive players (#1 and #2) are called the two-guard position used in a 5-on-5 game. Additionally, the start position of each offensive player is assigned to maintain a certain distance for observing the coordination process that the role of intervention decision and adjustment involves in the other players. On the defensive team, #1 and #2 confront offensive #1 and #2, and all three defenders’ starting positions are voluntary. The above information will help accurate player identification. Written informed consent was obtained from all the participants according to [32].
@@
Outdoor dataset. Outdoor dataset was constructed as a fixed-camera 3x3 basketball videos consisting of 12 videos totaling 143,276 frames for six players on the court. Fourteen male university basketball players, who regularly participate in official 5-on-5 games, created four teams (three players for each team; the remaining two players were reserved for accidents). A total of 12 games were played following a double round-robin format (i.e., each team participated in 6 games), and the four teams wore bibs in distinct colors: green, purple, pink, and yellow. The measurement was conducted in a university outdoor 3x3 court using two smartphone cameras (iPhone 13: $3840\times 2160$); but in this dataset we used only one camera. The 3x3 game was played according to the official 3x3 rules [^1] FIBA 3x3 official rules https://fiba3x3.com/en/rules.html except that the game was limited to 5 minutes without timeouts and without a knockout rule (21 points). This measurement was approved by the ethical committee of the Nagoya University and Ryutsu Keizai University and written informed consent was obtained from all the participants.
@@
Drone dataset. Drone dataset was constructed from a drone camera 3x3 basketball videos consisting of 6 videos (92 min. in total) but we extracted 4,999 frames for six players on the court due to the annotation issue. Sixteen male university basketball players, who regularly participate in official 5-on-5 games, created four teams (three players for each team; the remaining four players were reserved for accidents). A total of 6 games were played following a round-robin format (i.e., each team participated in 3 games). Since only black and white colors were available, one team wore black bibs while the other team wore white bibs to easily distinguish between the teams. The measurement was conducted in a university outdoor 3x3 court using two drone cameras (DJI Air 2S: $3840\times 2160$); but in this dataset, we used only one drone camera. The 3x3 game was played according to the official 3x3 rules1. This measurement was approved by the ethical committee of Anhui Normal University, China, and written informed consent was obtained from all the participants.
3.2 Annotation
@@
Field detection. In the Indoor dataset, field information was provided by [32], which was manually annotated based on the basketball court keypoints. Annotated keypoints in the center of a 5-centimeter wide court line, the court size was $9.50\times 15.05$ in the Indoor dataset and $11.05\times 15.05$ in the Outdoor dataset (depth $\times$ width [m], width corresponds to the length of the end line). Similarly, we manually annotated the field information in Outdoor dataset. However, in Drone dataset, although the field of view was generally stable, slight movements occurred due to wind and other environmental influences, and thus ground truth field information was not annotated. While field detection is a crucial component of the overall GSR task [70] (in particular, significantly impacting the holistic metric called GS-HOTA), this study focuses solely on the Track ID task using fixed-camera data. Consequently, only the Indoor and Outdoor datasets include field information, whereas the Drone dataset is used exclusively for tracking.
@@
Bounding boxes. In the Indoor dataset, bounding boxes (bboxes) were provided by [32]. In Outdoor and Drone datasets, we also manually annotated bboxes about each player. For all datasets, unique IDs for each player were assigned to enable tracking and identification over time.
@@
Figure 2: Our Track-ID baseline algorithm. Note that the court localization is performed manually and others can be automatically done.
4 TrackID3x3 Baseline Algorithms
@@
Here, we introduce the TrackID3x3 baseline algorithms employed to evaluate the performance of tracking, identification, and pose estimation methods on our dataset. First, we introduce a new Track-ID task, and its baseline algorithm, as well as tracking and 2D pose estimation baselines.
4.1 Track-ID Task
@@
One of our experimental tasks is a Track-ID task to obtain the location and identification information of players on the court from fixed-camera videos. It is similar to the recent GSR task [70], but the most significant difference is that the camera is fixed and shows the entire court. Our task is relatively easy and accurate to carry out court localization by manually annotating keypoints on a single image in the video. Another important difference is that referees are not the object of this task; in 3x3 basketball, since the two referees only move back and forth in their respective areas, their movements have little effect on the state of the game, and there is less incentive to track them than in soccer. Since only on-court players are targeted, the maximum number of detections required in each frame is determined to be 6.
@@
Note that there is a crucial difference between the Indoor and Outdoor datasets, in terms of player identification: the presence or absence of players’ jersey numbers. This is because the numbers were not printed on the bibs of the Indoor dataset, and this difference also accounts for the difference in sub-tasks performed between the two datasets. In addition, for the Outdoor dataset, the player is uniquely determined by two pieces of information: team affiliation and jersey number, because players on the same team wear the same color bibs with different numbers. Hence, the task formulation is also different in the two datasets.
@@
We formalize the Track-ID task according to the GSR task as follows. For each on-court player detection $d_{i}^{t}$ ($i\in\{1,2,...,6\}$) at frame $t$ in a video with a total number of frames $T$, predict its location and identification information. The information to be predicted in the Indoor dataset is,
@@
$d_{i}^{t}=\{\underbrace{\text{{court\_x, court\_y}}}_{\text{localization}},\;\underbrace{\text{{team, initial\_position}}}_{\text{identification}}\},$ (1)
@@
$d_{i}^{t}=\{\underbrace{\text{{court\_x, court\_y}}}_{\text{localization}},\;\underbrace{\text{{team, jersey\_number}}}_{\text{identification}}\},$ (2)
Track-ID Metric. The Track-ID task is evaluated by a metric called Track-ID HOTA (TI-HOTA). This is substantially equivalent to GS-HOTA but is given a different name to clarify the difference in the tasks performed. GS-HOTA is a metric that improves the similarity score for matching predictions and ground-truth from Higher Order Tracking Accuracy (HOTA) [50] to evaluate GSR, calculated only for pairs of predictions and ground-truth that are misaligned by less than the distance tolerance parameter $\tau\,[m]$ and have completely matching identification. Although the formulas are exactly the same, the difference between TI-HOTA and GS-HOTA is in the subtasks: manual or automatic court localization and the different identification information required.
Therefore, TI-HOTA is calculated as follows:
@@
where
$\operatorname{DetA}_{\alpha}=\frac{|TP|}{|TP|+|FP|+|FN|}$ (4)
$\operatorname{AssA}_{\alpha}=\frac{1}{|TP|}\sum_{c\in\{{TP}\}}\mathcal{A}(c)$ (5)
$\mathcal{A}(c)=\frac{|TPA(c)|}{|TPA(c)|+|FPA(c)|+|FNA(c)|}.$ (6)
$|TP|,|FP|$, and $|FN|$ are the number of true positives, false positives, and false negatives at threshold $\alpha$ of similarity score as described below, and $|TPA(c)|,|FPA(c)|$ and $|FNA(c)|$ are the number of true positive associations, false positive associations, and false negative associations. Note that $c$ is not a track based on a target ID in MOT, but a track defined by two identification information defined in equations (1, 2). The similarity score used for matching predictions and ground truth is as follows:
$\operatorname{Sim}_{\mathrm{TI\text{-}HOTA}}(P,G)=\operatorname{LocSim}(P,G)\times\operatorname{IdSim}(P,G),$ (7)
$\operatorname{LocSim}(P,G)=e^{\ln(0.05)\frac{\|P-G\|_{2}^{2}}{\tau^{2}}},$ (8)
$\operatorname{IdSim}(P,G)=\begin{cases}1&\text{ if all attributesmatch, }\\ 0&\text{ otherwise. }\end{cases}$ (9)
where $P$, $G$ are prediction and ground truth.
In addition to the TI-HOTA, this paper reports its sub-scores ($\operatorname{DetA}_{\alpha},\operatorname{AssA}_{\alpha},|FP|$, and $|FN|$) for detail evalutation.
Dataset Length [s] TI-HOTA (%)↑ DetA↑ AssA↑ FN↓ FP↓
Indoor 8.98 $\pm$ 3.24 80.75 $\pm$ 13.16 79.46 $\pm$ 14.42 82.11 $\pm$ 11.88 155.81 $\pm$ 140.10 133.81 $\pm$ 135.11
Outdoor 40.11 $\pm$ 34.23 46.11 $\pm$ 20.55 42.94 $\pm$ 20.87 49.81 $\pm$ 20.54 3695.74 $\pm$ 4280.34 3558.19 $\pm$ 4154.12
Table 3: Experimental results of the Track-ID task for the Indoor and the Outdoor dataset. Note that the mean ± SD at $\tau=0.5$ is presented.
Method Length [s] HOTA(%)↑ FN↓ FP↓ IDs↓
ByteTrack [95] 41.66 $\pm$ 4.44 47.92 $\pm$ 5.98 211.25 $\pm$ 93.93 28.25 $\pm$ 16.74 19.75 $\pm$ 4.92
BoT-SORT-ReID [1] 41.66 $\pm$ 4.44 50.64 $\pm$ 3.03 214.75 $\pm$ 110.45 19.5 $\pm$ 10.63 15.5 $\pm$ 1.91
Table 4: ByteTrack and BoT-SORT results on Drone dataset, IDs denote ID switches.
4.2 Track-ID Baseline Algorithm
People Detection and Tracking. We use BoT-SORT-ReID [1] with a pre-trained YOLOX [24] detecter without fine-tuning to track people in the videos. It is one of the strong image-based MOT methods in 5-on-5 basketball [28].
On-Court Player Classification. Since the images include people who are playing and those who are not, it is necessary to distinguish them. We employ a rule-based heuristic method based on court position coordinates. For the Indoor dataset, in each frame, detections are filtered to those present in the court when the virtual court line parallel to the end line is moved one-meter away perpendicular to the end line, and if more than seven detections remain, the number is reduced to six by excluding those closest to the end line. For the Outdoor dataset, we exclude detections that never stayed in the court for more than 10 consecutive frames. Although this method often leaves seven or more detections without eliminating the referees, the subsequent process narrows the number of detections down to six.
@@ 4.2 Track-ID Baseline Algorithm
Player Appearance Information Extraction. As mentioned earlier, there is a difference between Indoor and Outdoor in terms of the presence or absence of jersey numbers. For the Indoor, the color histogram of the player image regions segmented by Dectron2 [86] was calculated for each tracklet. Specifically, people segmentation was first performed for each frame, and the temporal median of the 3D color histograms ($8\times 8\times 8$ bin) was calculated for the segmentation region with the largest area in the bboxes of each player’s tracklet. For the Outdoor, a tracklet-level jersey number recognition pipeline [41] was used to predict the jersey number and calculate the color histogram of each tracklet using the intermediate product, the torso region images. This color histogram is similar in shape to Indoor and is the temporal median. The color histogram is the median value over up to 100 frames from the frame number with the smallest frame number, and its shape is the same as the Indoor. In both approaches, color histograms were used because of their lower computational cost than deep learning-based ReID features.
Tracklet Integration and Attributes Identification. For the Indoor, the first step was the integration of tracklets: each frame after the first was scanned in chronological order, and if only one new tracklet appeared in a frame, it was merged with the missing tracklet ID; if two or more new tracklets appeared simultaneously, similar ones were merged based on the similarity of their 3D color histograms. Note that Jensen-Shannon divergence was used as the similarity measure.
@@
Not only does the Indoor dataset not have information on the player’s jersey number, but all the players wear different colored bibs. Hence, the method of identifying the players’ team affiliation is limited to determining it practically from her relative position at the beginning of the video. Specifically, in the opening frame of the video, three pairs were created by pairing players detected at close distances to each other, with the one closer to the midpoint of the end line as the defense and the one farther away as the offense. The top, left, and right were determined relative to the midpoint of each pair’s position.
@@
As for the Outdoor, it is more complicated: first, exclude tracklets whose jersey numbers could not be recognized. Next, the tracklet closest to a certain position in the first frame is the offense: the midpoint of the court line parallel to the end line in the case of videos where the offense starts at the top, or the midpoint of the free throw line in the case of videos where the offense begins with a free throw. Then, in the case of a video with the offense starting at the top, the tracklet closest to the determined offensive one is the defense, and then the two tracklets with the highest color histgram similarity to the offensive one are the offense and the rest are the defense. In the case where the video begins with a free throw, the two tracklets with the most similar color histograms to the first determined offensive tracklet shall be the offense and the remainder shall be the defense. For subsequent frames, if new tracklets appear, they will be assigned to the team with more player shortages (or defense if there is a tie). Finally, if the number of team members is greater than three in each frame, the tracklets are removed in order of length, starting with the shortest.
4.3 Tracking Baselines in Drone Dataset
In Drone dataset, we describe the tracking baselines. We extracted four sequences from a complete 3v3 basketball game video, capturing critical game moments from the initial ball possession to termination events such as fouls, violations, or out-of-bounds plays. The dataset comprises 4,999 frames across these four sequences, providing a comprehensive representation of dynamic player movements in competitive gameplay scenarios. We implement two conventional baseline MOT algorithms: ByteTrack [95] and BoT-SORT-ReID [1] to track player movements.
For performance evaluation, we use HOTA [50] metric due to its comprehensive ability to measure both detection and association performance. Additionally, we incorporate supplementary metrics including FN, FP, and ID switches to provide a more granular analysis of tracking performance.
4.4 Pose Estimation Baselines

### v2

@@ 3. TrackID3x3 Dataset
As described in Table 2, we propose TrackID3x3 dataset composed of three subsets: Indoor, Outdoor, and Drone datasets (the novelty of our dataset is described in Introduction and Table 1. In this section, we describe the data collection and annotations in this order.
3.1. Dataset collection
@@
Indoor dataset. As described in (Ichikawa et al., 2024), Indoor dataset was constructed a fixed-camera 3x3 basketball videos consisting of 42 videos totaling 7,531 frames for six players on the court. The six female university basketball players regularly participate in official 5-on-5 games. The measurement was conducted in a university gymnasium (half-court of 5-on-5 basketball court) using a Sony HDR-CX680 camera ($1280\times 720$). Thus, only this subset has lower resolution compared to other subsets, with a lot of motion blur, making it relatively difficult to detect players in some scenes. For the mini-game design, offensive and defensive teams were fixed with players being assigned to specific roles. The ball handler of #1 is fixed, and the player begins without directly passing to #3. The start positions of the two offensive players (#1 and #2) are called the two-guard position used in a 5-on-5 game. Additionally, the start position of each offensive player is assigned to maintain a certain distance for observing the coordination process that the role of intervention decision and adjustment involves in the other players. On the defensive team, #1 and #2 confront offensive #1 and #2, and all three defenders’ starting positions are voluntary. The above information will help with accurate player identification. Written informed consent was obtained from all the participants according to (Ichikawa et al., 2024).
@@
Outdoor dataset. Outdoor dataset was constructed as a fixed-camera 3x3 basketball videos consisting of 12 videos totaling 143,276 frames for six players on the court. Fourteen male university basketball players, who regularly participate in official 5-on-5 games, created four teams (three players for each team; the remaining two players were reserved for accidents). A total of 12 games were played following a double round-robin format (i.e., each team participated in 6 games), and the four teams wore bibs in distinct colors: green, purple, pink, and yellow. The measurement was conducted in a university outdoor 3x3 court using two smartphone cameras (iPhone 13: $3840\times 2160$); but in this dataset we used only one camera. The 3x3 game was played according to the official 3x3 rules [^1] FIBA 3x3 official rules https://fiba3x3.com/en/rules.html except that the game was limited to 5 minutes without timeouts and without a knockout rule (21 points). This measurement was approved by the ethical committee of Nagoya University and Ryutsu Keizai University and written informed consent was obtained from all the participants.
@@
Drone dataset. Drone dataset was constructed from a drone camera 3x3 basketball videos consisting of 6 videos (92 min. in total) but excluding out-of-play sections, 127,119 frames were extracted focusing on the six players on the court. Sixteen male university basketball players, who regularly participate in official 5-on-5 games, created four teams (three players for each team; the remaining four players were reserved for accidents). A total of 6 games were played following a round-robin format (i.e., each team participated in 3 games). Since only black, brown and white colors were available, one team wore black or brown jersey while the other team wore white jersey to easily distinguish between the teams. The measurement was conducted in a university outdoor 3x3 court using two drone cameras (DJI Air 2S: $3840\times 2160$), but in this dataset, we used only one drone camera. There are some camera movements, but it is not as dramatic as in broadcast footage in existing benchmark datasets, and the players are in the camera’s field of view in almost every frame. The 3x3 game was played according to the official 3x3 rules1. This measurement was approved by the ethical committee of Anhui Normal University, China, and written informed consent was obtained from all the participants.
3.2. Annotation
@@
Field detection. In the Indoor dataset, field information was provided by (Ichikawa et al., 2024), which was manually annotated based on the basketball court keypoints. Annotated keypoints in the center of a 5-centimeter wide court line, the court size was $9.50\times 15.05$ in the Indoor dataset and $11.05\times 15.05$ in the Outdoor dataset (depth $\times$ width [m], width corresponds to the length of the end line). Similarly, we manually annotated the field information in Outdoor dataset. However, in Drone dataset, although the field of view was generally stable, slight movements occurred due to wind and other environmental influences, and thus ground truth field information was not annotated. While field detection is a crucial component of the overall GSR task (Somers et al., 2024b) (in particular, significantly impacting the holistic metric called GS-HOTA), this study focuses solely on the Track-ID task using fixed-camera data. Consequently, only the Indoor and Outdoor datasets include field information, whereas the Drone dataset is used exclusively for tracking.
@@
Bounding boxes. In the Indoor dataset, bounding boxes (bboxes) were provided by (Ichikawa et al., 2024). In the Outdoor dataset and Drone dataset, approximately half of the data was annotated manually, and the remaining half was annotated using a method adopted in a previous study (Xu et al., 2024), in which the output of the tracking algorithm was manually corrected (BoT-SORT-ReID (Aharon et al., 2022) was used as the tracking algorithm in this study). For all datasets, unique IDs for each player were assigned to enable tracking and identification over time.
@@
Figure 2. Our Track-ID baseline algorithm. Note that the court localization is performed manually and others can be automatically done.
4. TrackID3x3 Baseline Algorithms
@@
Here, we introduce the TrackID3x3 baseline algorithms employed to evaluate the performance of tracking, identification, and pose estimation methods on our dataset. First, we introduce a new Track-ID task, and its baseline algorithm, as well as tracking (Appendix A) and 2D pose estimation baselines.
4.1. Track-ID Task
@@
One of our experimental tasks is a Track-ID task to obtain the location and identification information of players on the court from fixed-camera videos. It is similar to the recent GSR task (Somers et al., 2024b), but the most significant difference is that the camera is fixed and shows the entire court. Our task is relatively easy and accurate to carry out court localization by manually annotating keypoints on a single image in the video. Another important difference is that referees are not the object of this task; in 3x3 basketball, since the two referees only move back and forth in their respective areas, their movements have little effect on the state of the game, and there is less incentive to track them than in soccer. Since only on-court players are targeted, the maximum number of detections required in each frame is determined to be 6.
@@
Note that there is a crucial difference between the Indoor and Outdoor datasets, in terms of player identification: the presence or absence of players’ bib numbers. This is because the numbers were not printed on the bibs of the Indoor dataset, and this difference also accounts for the difference in sub-tasks performed between the two datasets. In addition, for the Outdoor dataset, the player is uniquely determined by two pieces of information: team affiliation and jersey number, because players on the same team wear the same color bibs with different numbers. Hence, the task formulation is also different in the two datasets.
@@
We formalize the Track-ID task in 3x3 basketball according to the GSR task as follows. For each on-court player detection $d_{i}^{t}$ ($i\in\{1,2,...,6\}$) at frame $t$ in a video with a total number of frames $T$, predict its location and identification information. The information to be predicted in the Indoor dataset is,
@@
(1) $d_{i}^{t}=\{\underbrace{\text{{court\_x, court\_y}}}_{\text{localization}},\;\underbrace{\text{{team, initial\_position}}}_{\text{identification}}\},$
where court_x and court_y are the $x,y$ coordinates in the court coordinate system, team $\in\{\text{offense, defense}\}$ indicates whether the player was on the offensive or defensive team at the start of the video, and initial_position $\in\{\text{top, left, right}\}$ indicates the position at the start of the video. For each attribute of the ground truth initial_position, the top is given to two players who start the offense and defend her by the top of the key (no check-ball is played and the former holds the ball from the beginning of the video). The left and right are given relative to the coordinates of the other players. The information that should be predicted in the Outdoor dataset is,
(2) $d_{i}^{t}=\{\underbrace{\text{{court\_x, court\_y}}}_{\text{localization}},\;\underbrace{\text{{team, jersey\_number}}}_{\text{identification}}\},$
Track-ID Metric. The Track-ID task is evaluated by a metric called Track-ID HOTA (TI-HOTA). This is essentially equivalent to GS-HOTA but is given a different name to clarify the difference in the tasks performed. GS-HOTA is a metric that improves the similarity score for matching predictions and ground-truth from Higher Order Tracking Accuracy (HOTA) (Luiten et al., 2021) to evaluate GSR, calculated only for pairs of predictions and ground-truth that are misaligned by less than the distance tolerance parameter $\tau\,[m]$ and have completely matching identification. Although the formulas are exactly the same, the difference between TI-HOTA and GS-HOTA is in the subtasks: manual or automatic court localization and the different identification information required.
Therefore, TI-HOTA is calculated as follows:
(3) $\text{TI-HOTA}=\frac{1}{19}\sum_{\alpha\in\{0.05,0.10,\dots,0.95\}}\sqrt{\operatorname{TI-DetA}_{\alpha}\times\operatorname{TI-AssA}_{\alpha}},\\$
where
(4) $\operatorname{TI-DetA}_{\alpha}=\frac{|TP_{\alpha}|}{|TP_{\alpha}|+|FP_{\alpha}|+|FN_{\alpha}|},$
(5) $\operatorname{TI-AssA}_{\alpha}=\frac{1}{|TP_{\alpha}|}\sum_{c\in\{{TP_{\alpha}}\}}\mathcal{A}_{\alpha}(c),$
(6) $\mathcal{A}_{\alpha}(c)=\frac{|TPA_{\alpha}(c)|}{|TPA_{\alpha}(c)|+|FPA_{\alpha}(c)|+|FNA_{\alpha}(c)|}.\$
$|TP_{\alpha}|,|FP_{\alpha}|$, and $|FN_{\alpha}|$ are the number of true positives, false positives, and false negatives at threshold $\alpha$ of similarity score as described below, and $|TPA_{\alpha}(c)|,|FPA_{\alpha}(c)|$ and $|FNA_{\alpha}(c)|$ are the number of true positive associations, false positive associations, and false negative associations. Note that $c$ is a true positive pair that matched at least once in all frames based on the similarity score:
(7) $\operatorname{Sim}_{\mathrm{TI\text{-}HOTA}}(P,G)=\operatorname{LocSim}(P,G)\times\operatorname{IdSim}(P,G),$
(8) $\operatorname{LocSim}(P,G)=e^{\ln(0.05)\frac{\|P-G\|_{2}^{2}}{\tau^{2}}},$
@@
where $P$, $G$ are prediction and ground truth.
In addition to the TI-HOTA, this paper reports its sub-scores ($\operatorname{TI-DetA}$ and $\operatorname{TI-AssA}$) for detail evalutation.
Dataset Length [s] TI-HOTA (%)↑ TI-DetA (%)↑ TI-AssA (%)↑
Indoor 8.98 $\pm$ 3.24 85.53 $\pm$ 7.09 84.86 $\pm$ 8.01 86.23 $\pm$ 6.15
Outdoor 40.11 $\pm$ 34.23 71.03 $\pm$ 17.79 67.15 $\pm$ 20.12 75.51 $\pm$ 15.65
Dataset Length [s] TI-HOTA (%)↑ TI-DetA (%)↑ TI-AssA (%)↑
Indoor 8.98 $\pm$ 3.24 84.45 $\pm$ 7.44 83.60 $\pm$ 8.42 85.33 $\pm$ 6.44
Outdoor 40.11 $\pm$ 34.23 66.20 $\pm$ 18.07 61.97 $\pm$ 19.86 71.08 $\pm$ 16.42
Table 3. Experimental results of the Track-ID task; Left: w/ ID Switch Merging, Right: w/o ID Switch Merging. Note that the mean ± SD at $\tau=1$ is presented.
4.2. Track-ID Baseline Algorithm
People Detection and Tracking. We use CAMELTrack (Somers et al., 2025) with a pre-trained YOLOX (Ge et al., 2021) detecter without fine-tuning to track people in the videos. This tracking algorithm is one of the few methods that has reported a HOTA score of over 80% on the SportsMOT test set (Cui et al., 2023), and was selected for its high accuracy and code availability. The weights of the association module were pre-trained using SportsMOT (Cui et al., 2023). The weights used in YOLOX are the same as those used during annotation, but the post-processing settings were changed to reduce detection power. This is partly because it is the same detector used for annotation, so as to avoid using exactly the same detection results, but also because in basketball player tracking, it may be effective to use a detecter that does not have overly strong detection power (the detail explanation is described in Section 6).
Exclusion of Non-Players. Since the images include people who are playing and those who are not, it is necessary to distinguish them. We employ a rule-based heuristic method based on court position coordinates. Exclusion criteria are divided into detection level and tracklet level. There is one exclusion criterion for detection level: detections more than 3 meter away from the outer lines of the court are excluded. On the other hand, tracklet level exclusion is performed based on three criteria: tracks that do not exist continuously for 10 or more frames on the court, tracks that exist in more than half of the frames in the area along the end line, from 3 meters outside the court to 1 meter inside the court, and tracks that exist in more than half of the frames outside the extension lines of the two long sides of the paint area and more than 10 meter away from the end line (the area near the so-called coffin corner). The latter two criteria are intended to exclude referees. For datasets with jersey numbers, referees can be excluded using the results of jersey number recognition described later, but to lighten the processing load as much as possible, we use simple rules to exclude them here.
ID Switch Detection & Merging. After narrowing down the player’s tracklets, apply the module that detects and merges the ID Switches. The purpose of this module is to merge fragmented tracklets for the same player. Specifically, a tracklet ID that appeared midway through a video clip is recognized as an ID switch and merged with a tracklet ID that existed from the beginning if the simultaneous existence of those IDs frame count is below the threshold (for details, refer to Algorithm 1). Although it cannot handle ID exchanges between players (described later) and requires all players to be constantly visible on camera, it has fewer parameters to adjust than Global Tracklet Association (GTA) (Sun et al., 2024), which includes two clustering algorithms, and does not require the calculation of ReID features, resulting in lower computational costs. In this study, the only parameter to be adjusted, $T_{\text{overlap}}$, was set to 10 for Indoor and 50 for Outdoor, taking into account the average length of the video clips.
Algorithm 1 ID Switch Detection & Merging
1:DataFrame $D$ containing tracklets, overlap threshold $T_{\text{overlap}}$, total frame count of video clip $F_{\max}$
@@ 4.2. Track-ID Baseline Algorithm
[…]
6:for all newID in newIDs do
7:  $\text{candidates}\leftarrow[]$
8:  $\text{framesNew}\leftarrow\{\,f\mid D.\text{tracklet\_id}=\text{newID}\land\text{on\_court}(f)\}$
9:  for all origID in origIDs do
10:    $\text{framesOrig}\leftarrow\{\,f\mid D.\text{tracklet\_id}=\text{origID}\land\text{on\_court}(f)\}$
11:    $\text{overlap}\leftarrow|\text{framesOrig}\cap\text{framesNew}|$
12:    if $\text{overlap}\geq T_{\text{overlap}}$ then
13:     continue
14:    end if
15:    $\text{unionFrames}\leftarrow\text{framesOrig}\cup\text{framesNew}$
16:    $\text{missing}\leftarrow F_{\max}-|\text{unionFrames}|$
17:    $\text{cost}\leftarrow\text{overlap}+\text{missing}$
18:    Append $(\text{origID},\text{cost})$ to candidates
19:  end for
20:  if candidates is empty then
21:    continue
22:  end if
23:  $\text{bestID}\leftarrow$ origID in candidates with minimum cost
24:  Relabel all rows where tracklet_id = newID to tracklet_id = bestID
25:  Drop duplicate rows on (frame_id, tracklet_id), keeping the first
26:end for
@@
Jersey Number Recognition. As mentioned earlier, there is a difference between Indoor and Outdoor in terms of the presence or absence of bibs numbers. Therefore, only Outdoor requires recognition of jersey numbers. We used a trucklet-level jersey number recognition pipeline (Koshkina and Elder, 2024) to obtain one jersey number prediction result for each merged player tracklet. At the same time, we calculated the 3D color histogram for each tracklet using the cutout image of the torso region, which is an intermediate product. At the same time, we calculated the 3D color histogram for each tracklet using the cropped images of the torso region, which is an intermediate product. The color histogram has 8 × 8 × 8 bin and the median value over up to 100 frames from the smallest frame number.
@@
Attribute Identification. Not only does the Indoor dataset not have information on the player’s jersey number, but all the players wear different colored bibs. Hence, the method of identifying the players’ team affiliation is practically limited to determining it from her relative position at the beginning of the video. Specifically, in the opening frame of the video, three pairs were created by pairing players detected at close distances to each other, with the one closer to the midpoint of the end line as the defense and the one farther away as the offense. The top, left, and right were determined relative to the midpoint of each pair’s position.
@@
As for the Outdoor, it is more complicated: first, exclude tracklets whose jersey numbers could not be recognized to colpletely e referees. Next, the tracklet closest to a certain position in the first frame is the offense: the midpoint of the outer court line parallel to the end line in the case of videos where the offense starts at the top, or the midpoint of the free throw line in the case of videos where the offense begins with a free throw. Then, in the case of a video with the offense starting at the top, the tracklet closest to the determined offensive one is the defense, and then the two tracklets with the highest color histgram similarity based on Jensen-Shannon divergence to the offensive one are the offense and all the rest are the defense. In the case where the video begins with a free throw, the two tracklets with the most similar color histograms to the first determined offensive shall be the offense and all the remainder shall be the defensive.
Post-Processing. Finally, perform post-processing to produce the final output. Here, linear interpolation, end point extrapolation, and frame level detection number limitation are applied. Linear interpolation simply fills in the gaps in the tracklet linearly. Endpoint extrapolation fills in missing tracklets at the beginning or end of a video based on the difference between the two consecutive detections immediately after or before. The frame level detection limitation excludes detections in more than 6 detections, starting with those track with the fewest on-court frames, to ensure that there are 6 players or fewer.
4.3. Pose Estimation Baselines